OpenStudy (anonymous):

solve the system of equations. Enter your answer as an ordered pair , 4x-2y=-2 & 4x+5y=33

OpenStudy (anonymous):

There are many ways of doing this. The easiest for me is usually substitution. Isolate a variable in the first equation and then plug it into the second. In this case from the first equation we can gather y=1+2x. Plug this into the second equation and you get 4x+5(1+2x)=33. Solve this for x to get x=2. Now plug this value into your expression for y from the first equation (y=1+2x) to get y = 5. You can check your answer by plugging these values into the original system. 4(2)-2(5) = -2, 4(2)+5(5) = 33. Your answer is (2,5)